Solve for X, and just for real numbers
log(X+4)+log10(X-5)=1

Answer by stanbon(75887)

Solve for X, and just for real numbers
log(X+4)+log10(X-5)=1
----
log[(x+4)(x-5)] = 1
---
log[x^2-x-20] = 1
---
x^2-x-20 = 10
----
x^2-x-30 = 0
Factor:
(x-6)(x+5) = 0
x = 6 or x = -5
===================
Cheers,
Stan H.
====================